OpenStudy (anonymous):

help me out with this problem please, i got 8=16, wolfram got some decimal number, but the answer key i have has -2

OpenStudy (anonymous):

\[\sqrt{x+6}+\sqrt{2-x}=4\] help please

OpenStudy (anonymous):

You got no choice but to square both sides of this guy, lets do it toghether

OpenStudy (anonymous):

i was told to use the power rule which means to deal with one radical sign at a time, i don't get it. i tried squaring both sides the way it is but i got a wrong answer

OpenStudy (anonymous):

what exactly are we doing here?

OpenStudy (anonymous):

i'm supposed to isolate a radical sign on one side of the equation that would give me \[\sqrt{x+6}=4-\sqrt{2-x}\] and then I should square both sides, i get stuck when i get to squaring what's on the right of the equal sign

OpenStudy (anonymous):

okay, you should get: